(Rana, Suresh, Jenne, Kiwamu, Kevin, Yuta)
Last night we locked MC by feeding back the signal to NPRO PZT.
But it was not so stable.
We wanted more gain to lower the seismic motion, but we don't need high gain in high frequency part(>~1kHz) because it may cause something bad(NRPO PZT oscilatting, PMC not able to catch up with the NPRO frequency change, etc).
So, we put DC gain boost today.
It successfully made MC locking stable!
What we did:
1. Lowered the main laser temperature from 32.2° C to 31.8° C.
When we increased the laser temperature, PMC transmission get lower. 32.2° C was on the cliff, so we put it to plateau region.
2. Lowered the gain of PMC servo (2dB instead of 8dB last night), because PMC was oscillating.
We got 5.3V OMC transmission.
3. Made 3Hz pole, 30Hz zero filter and put in NPRO PZT servo loop.
0dB at DC, -20dB at high frequency (see Kevin's elog #3802)
4. Put more gain to NPRO PZT servo loop to compensate -20dB at high frequency.
In a word, we put DC gain boost.
Attachment #1 is the MEDM screen screenshot for MC servo.
5. Aligned the beam into QPD at MC2 trans.
We put lens in front of the QPD.
Now, we can see the actual motion of the beam, and resonance peaks(Attachment #2; not locked at the highest).
(We added 30Hz LPF after each 4 quadrant inputs to reduce noise)
- optimize MC suspension alignments
- activate OL DAQ channels
- reduce RFAM
- install tri-mod box
- QPD signal at MC2 should be more high(currently, ~7counts which equals to ~4mV)
- change temperature of X-end laser to get green beat